The perimeter of a rectangle is 154 inches. The length exceeds the width by 39 inches. Find the length and width.
Answer by checkley71(8403)

L=W+39
2L+2W=154
2(W+39)+2W=154
2W+78+2W=154
4W=154-78
4W=76
W=76/4
W=19 FOR THE WIDTH
L=19+39=58 FOR THE LENGTH
PROOF
2*58+2*19
116+38=154
154=154